# UWPD: A General Paradigm for Invisible Watermark Detection Agnostic to Embedding Algorithms

### 06 Mar 2026 v1

@@ 5.3 Frequency Shield Decoder
To ensure high decoupling of the network architecture and universal applicability to different backbones (such as CNNs and Transformers), our decoder adopts a lightweight design. The deep 2D feature map $\bm{Z}\in\mathbb{R}^{C\times H^{\prime}\times W^{\prime}}$ outputted by DMSA is compressed into a 1D vector $\bm{z}\in\mathbb{R}^{C}$ through global adaptive average pooling. Subsequently, it is mapped to binary classification logits through a two-layer MLP with Dropout:
$\hat{\bm{y}}=\bm{W}_{2}\big(\text{ReLU}(\text{BatchNorm}(\bm{W}_{1}\bm{z}))\big)$ (12)
where ReLU is the ReLU activation function, and $\bm{W}_{1},\bm{W}_{2}$ are the fully connected layer weights. Since the complex frequency-domain feature mining has been undertaken by the front-end ASPM and DMSA modules, this streamlined classification head effectively prevents the model from overfitting when facing extremely weak watermark features.
6 Experiments
In the experimental phase, we evaluated various baseline models as well as our proposed FSNet model on UniFreq-100K, thereby verifying the rationality of the dataset and the effectiveness of the FSNet structure, and further providing baseline standards for the UWPD task. Meanwhile, we also conducted additional experiments to further explore the UWPD task and related models. All experiments adopt Cross-Entropy as the loss function and use the AdamW [17] optimizer for training. The learning rate is searched within the range {0.0001, 0.0005, 0.001, 0.005, 0.01}, and the batch size is fixed at 128. All experiments are completed on two NVIDIA A100 GPUs. We adopt Accuracy and F1-score as the core evaluation metrics.
6.1 Model Evaluation
The UniFreq-100K dataset contains 9 types of invisible watermarks. To evaluate the models’ generalization ability to recognize unknown watermarks on the UWPD task, we adopted a "Leave-One-Out" cross-evaluation strategy: during the training phase, the model learns the common features of 8 types of watermarks, and during the testing phase, it is used to identify the 9th unseen watermark. In terms of dataset splitting, for each evaluation, we combine the single type of watermark images serving as the test target with an equal number of unwatermarked images to form the test set, and combine the remaining 8 types of watermark images with an equal number of unwatermarked images to form the training set, thereby ensuring that the ratio of positive to negative samples is strictly maintained. The cross-evaluation results of each model on the 9 watermark types are detailed in Table 2.
Table 2: Comparison of different models on various watermarking and steganography methods (Accuracy and F1-score)
@@ 6.1 Model Evaluation
As seen from the experimental results in Table 2, in the vast majority of cases, the accuracy of all models exceeds 60%, proving that commonalities do indeed exist in invisible watermark embedding technologies, and universal recognition models can be constructed to identify the presence of invisible watermarks. Among them, due to its unique high-frequency capturing design, FSNet outperforms baseline models not specifically designed for this task in most cases, further demonstrating that enhancing high-frequency signals in the UWPD task is beneficial for improving model performance.
However, it is noteworthy that on LSB and Patchwork, the recognition accuracies of baseline models and FSNet are both lower than 60%. With a positive-to-negative sample ratio of 1:1, such accuracy indicates that the models have not achieved effective classification. We believe this is related to the fact that these two methods rely on bit features rather than directly embedding high-frequency signals. We will further discuss this issue in the Appendix.
6.2 Dataset Analysis
6.2.1 Scale
6.2.2 Relationships Between Similar Datasets
Among the 9 watermark types, there exist similar algorithms, such as DCT and DWT. Here we explore the influence between these similar algorithms. We conducted two experimental schemes: (1) We completely removed the DCT and DWT frequency-domain algorithms, trained on the remaining data, and then evaluated on DCT and DWT datasets. (2) We completely removed the deep learning-based HiDDeN and StegaStamp algorithms, trained on the remaining data, and evaluated on HiDDeN and StegaStamp datasets. The experimental results are shown in Figure 4.
The results in the table show that compared to scenarios where similar algorithms are not removed, model performance has declined, indicating that features among similar algorithms exert a certain influence on recognition. However, FSNet suffered less performance degradation due to its capability of capturing the inherent high-frequency features of invisible watermarks.
@@ 7 Conclusion
Aiming at the current issue of lacking universal detection methods for invisible watermarks, this study proposes the UWPD task and constructs a universal invisible watermark detection method. By constructing a dataset covering various invisible watermark algorithms, conditions are provided for universal model training. Targeting the common high-frequency features of invisible watermarks, we designed the FSNet model, which can effectively capture the invisible watermark information hidden in high frequencies, achieving optimal performance on the UWPD task and providing modeling ideas and a baseline framework for future research.
8 Future Work
The UWPD task is currently still in its infancy. The detection accuracy for bit-based algorithms such as LSB in the constructed dataset is relatively low, requiring further research in the future. Meanwhile, starting from around 30% of the dataset, the performance improvement of the model is relatively limited, yet there is still considerable room for improvement in the detection accuracy of models on the UniFreq-100K dataset. Enabling the models to better capture the commonalities of invisible watermarks will be the focus of subsequent research. In addition, the models’ performance varies greatly across different data. For example, Swin Transformer achieves an accuracy of 91.5% on the DWT dataset, far exceeding other baselines, while FSNet also underperforms compared to other baseline models on DCT and Stable Signature. This indicates that model architecture has a profound impact on the UWPD task, leaving massive room for future model design research.
Appendix 0.A UniFreq-100K Dataset Construction Details
To comprehensively evaluate the model’s performance on the Universal Watermark Presence Detection (UWPD) task, we constructed the large-scale UniFreq-100K dataset. This appendix elaborates on the sources of various watermark carrier images in the dataset, the embedding settings for different watermarking algorithms, specific engineering implementation details, and related copyright statements.
0.A.1 Selection and Sampling of Watermarked Images
@@
0.A.2 Watermarking Algorithm and Parameter Settings
@@
0.A.2.1 Traditional Spatial and Frequency Domain Baseline Algorithms
@@
0.A.2.2 Deep Learning Watermarking Algorithms
@@
0.A.2.3 Generative and AIGC Watermarking Algorithms
@@
0.A.3 Dataset Copyright and Open-Source Statement
@@
Appendix 0.B Degradation Analysis of Universal Invisible Watermark Detection Performance Based on Weak Spatial Characteristics
@@
0.B.1 Patchwork: Extreme Spatial Sparsity and Signal Annihilation During Downsampling
@@
0.B.2 LSB: Extremely Low-Amplitude Perturbations and Feature Masking at the Normalization Level
@@
0.B.3 Accurate Capture of Dense Frequency Anomalies and Universal Detection Boundaries
@@
The experiments and derivations above profoundly reveal the capability boundaries of current models in the Universal Watermark Presence Detection (UWPD) task. Existing visual models (including FSNet, which performs specific high-frequency enhancement) inherently rely heavily on capturing continuous frequency anomalies with a certain energy density. LSB and Patchwork successfully bypassed the feature extraction preferences of modern networks by exploiting extreme spatial sparsity and low amplitude. However, because these early algorithms completely lack robustness against routine image post-processing like JPEG compression, they have been eliminated in real-world copyright protection scenarios. Therefore, the models’ performance on such algorithms paradoxically proves their accurate locking capability on the core common feature of modern invisible watermarks: “dense high-frequency spectral anomalies.”
Appendix 0.C Visual Analysis of the Feature Capture Mechanism of FSNet
In the analyses of the main text and Appendix B, we theoretically demonstrated the dilemmas faced by traditional visual baseline models in the Universal Watermark Presence Detection (UWPD) task. To intuitively dissect how FSNet breaks through this limitation and achieves outstanding zero-shot generalization capabilities under extremely weak signals, we designed two sets of in-depth controlled visualization experiments targeting the network’s core innovative modules (ASPM and DMSA).
@@ Appendix 0.C Visual Analysis of the Feature Capture Mechanism of FSNet
0.C.1 Shallow Frequency Domain Gating Analysis
@@
0.C.2 Deep Multi-band Attention Analysis
Retaining high-frequency signals at shallow layers is not sufficient to handle the complex heterogeneity of unknown watermarking algorithms in the UWPD task. Modern invisible watermarks often concentrate energy in specific narrow-band frequencies. The primary design intention of the Dynamic Multi-band Attention mechanism (DMSA) is to recalibrate channel weights in the deep feature space using multi-branch discrete cosine transforms and extremum pooling.

### 13 Mar 2026 v2

@@ 5.3 Frequency Shield Decoder
To ensure high decoupling of the network architecture and agnostic applicability to different backbones (such as CNNs and Transformers), our decoder adopts a lightweight design. The deep 2D feature map $\bm{Z}\in\mathbb{R}^{C\times H^{\prime}\times W^{\prime}}$ outputted by DMSA is compressed into a 1D vector $\bm{z}\in\mathbb{R}^{C}$ through global adaptive average pooling. Subsequently, it is mapped to binary classification logits through a two-layer MLP with Dropout:
$\hat{\bm{y}}=\bm{W}_{2}\big(\text{ReLU}(\text{BatchNorm}(\bm{W}_{1}\bm{z}))\big),$ (12)
where ReLU is the ReLU activation function, and $\bm{W}_{1},\bm{W}_{2}$ are the fully connected layer weights. Since the complex frequency-domain feature mining is handled by the front-end ASPM and DMSA modules, this streamlined classification head effectively prevents the model from overfitting on extremely weak watermark features.
Table 1: Comparison of different models on various watermarking and steganography methods (Accuracy and F1-score). Bold: best performance; Italic: second-best performance.
@@
6 Experiments
In the experimental phase, we evaluated various baseline models as well as our proposed FSNet on UniFreq-100K, thereby verifying the validity of the dataset and the effectiveness of the FSNet architecture, and further providing baseline standards for the AWPD task. Meanwhile, we also conducted additional experiments to further explore the AWPD task and related models. All experiments adopt Cross-Entropy as the loss function and use the AdamW [17] optimizer for training. The learning rate is searched within the range {0.0001, 0.0005, 0.001, 0.005, 0.01}, and the batch size is fixed at 128. All experiments are completed on two NVIDIA A100 GPUs. We adopt Accuracy and F1-score as the core evaluation metrics.
6.1 Model Evaluation
The UniFreq-100K dataset contains 9 types of invisible watermarks. To evaluate the models’ generalization ability to recognize unknown watermarks on the AWPD task, we adopted a ”Leave-One-Out” cross-evaluation strategy: during the training phase, the model learns the common features of 8 types of watermarks, and during the testing phase, it is used to identify the 9th unseen watermark. In terms of dataset splitting, for each evaluation, we combine the single type of watermark images serving as the test target with an equal number of unwatermarked images to form the test set, and combine the remaining 8 types of watermark images with an equal number of unwatermarked images to form the training set, thereby ensuring that the ratio of positive to negative samples is strictly maintained. The cross-evaluation results of each model on the 9 watermark types are detailed in Table 1.
@@ 6.1 Model Evaluation
As seen from the experimental results in Table 1, in the vast majority of cases, the accuracy of all models exceeds 60%, proving that commonalities do indeed exist in invisible watermark embedding technologies, and agnostic recognition models can be constructed to identify the presence of invisible watermarks. Among them, due to its unique high-frequency capturing design, FSNet outperforms baseline models not specifically designed for this task in most cases, further demonstrating that enhancing high-frequency signals in the AWPD task is beneficial for improving model performance.
However, it is noteworthy that on LSB and Patchwork, the recognition accuracies of baseline models and FSNet are both lower than 60%. With a positive-to-negative sample ratio of 1:1, such accuracy indicates ineffective classification. We attribute this to the fact that these two methods rely on bit-level manipulation rather than directly embedding high-frequency signals. We will further discuss this issue in the Appendix.
6.2 Dataset Analysis
6.2.1 Scale
6.2.2 Relationships Between Similar Datasets
Among the 9 watermark types, certain algorithms share similarities, such as DCT and DWT. Here we explore the mutual influence of these similar algorithms. We conducted two experimental schemes: (1) We completely removed the DCT and DWT frequency-domain algorithms, trained on the remaining data, and then evaluated on DCT and DWT datasets. (2) We completely removed the deep learning-based HiDDeN and StegaStamp algorithms, trained on the remaining data, and evaluated on HiDDeN and StegaStamp datasets. The experimental results are shown in Figure 4.
The results indicate that model performance declines when similar algorithms are removed, suggesting that shared features among similar algorithms influence recognition. However, FSNet suffered less performance degradation due to its capability of capturing the inherent high-frequency features of invisible watermarks.
@@ 7 Conclusion
Aiming at the current issue of lacking agnostic detection methods for invisible watermarks, this study proposes the AWPD task and constructs an agnostic invisible watermark detection method. By constructing a dataset covering various invisible watermark algorithms, conditions are provided for agnostic model training.Targeting the common high-frequency features of invisible watermarks, we designed the FSNet model. It can effectively capture the invisible watermark information hidden in high frequencies, achieving optimal performance on the AWPD task and providing a baseline framework for future research.
Appendix A UniFreq-100K Dataset Construction Details
To comprehensively evaluate the model’s performance on the Agnostic Watermark Presence Detection (AWPD) task, we constructed the large-scale UniFreq-100K dataset. This appendix elaborates on the sources of various watermark carrier images in the dataset, the embedding settings for different watermarking algorithms, specific engineering implementation details, and related copyright statements.
A.1 Selection and Sampling of Watermarked Images
@@
A.2 Watermarking Algorithm and Parameter Settings
@@
A.2.1 Traditional Spatial and Frequency Domain Baseline Algorithms
@@
A.2.2 Deep Learning Watermarking Algorithms
@@
A.2.3 Generative and AIGC Watermarking Algorithms
@@
A.3 Detailed Cross-Distribution Table
Table 3 presents the complete cross-distribution of the UniFreq-100K dataset, showing the exact number of images for each combination of image category and watermarking algorithm.
Table 3: Detailed cross-distribution of image categories and watermarking algorithms in the UniFreq-100K dataset (all values in thousands)
Algorithm Real 2D/Art AIGC E-Post. Scan. Total
Unwatermarked 30 12 30 12 12 96
LSB 4 2 2 2 2 12
Patchwork 4 2 2 2 2 12
DCT 4 2 2 2 2 12
DWT 4 2 2 2 2 12
HiDDeN 4 2 2 2 2 12
StegaStamp 4 2 2 2 2 12
Stable Sig. 0 0 10 0 0 10
Tree-Ring 0 0 10 0 0 10
SynthID 0 0 2 0 0 2
Watermarked Sub. 24 12 34 12 12 94
Grand Total 54 24 64 24 24 190
As shown in Table 3, the dataset maintains a balanced distribution across different image categories while accommodating the specific characteristics of different watermarking algorithms. Notably, generative watermarking algorithms (Stable Signature, Tree-Ring, and SynthID) are exclusively applied to AIGC images, as these algorithms are specifically designed for integration with generative models.
A.4 Dataset Copyright and Open-Source Statement
@@
Appendix B Degradation Analysis of Agnostic Invisible Watermark Detection Performance Based on Weak Spatial Characteristics
@@
B.1 Patchwork: Extreme Spatial Sparsity and Signal Annihilation During Downsampling
@@
B.2 LSB: Extremely Low-Amplitude Perturbations and Feature Masking at the Normalization Level
@@
B.3 Accurate Capture of Dense Frequency Anomalies and Agnostic Detection Boundaries
@@
The experiments and derivations above profoundly reveal the capability boundaries of current models in the AWPD task. Existing visual models (including FSNet, which performs specific high-frequency enhancement) inherently rely heavily on capturing continuous frequency anomalies with a certain energy density. LSB and Patchwork successfully bypassed the feature extraction preferences of modern networks by exploiting extreme spatial sparsity and low amplitude. However, because these early algorithms completely lack robustness against routine image post-processing like JPEG compression, they have been eliminated in real-world copyright protection scenarios. Therefore, the models’ performance on such algorithms paradoxically proves their accurate locking capability on the core common feature of modern invisible watermarks: “dense high-frequency spectral anomalies.”
Appendix C Visual Analysis of the Feature Capture Mechanism of FSNet
In the analyses of the main text and Appendix B, we theoretically demonstrated the dilemmas faced by traditional visual baseline models in the AWPD task. To intuitively dissect how FSNet breaks through this limitation and achieves outstanding zero-shot generalization capabilities under extremely weak signals, we designed two sets of in-depth controlled visualization experiments targeting the network’s core innovative modules (ASPM and DMSA).
@@ Appendix C Visual Analysis of the Feature Capture Mechanism of FSNet
C.1 Shallow Frequency Domain Gating Analysis
@@
C.2 Deep Multi-band Attention Analysis
Retaining high-frequency signals at shallow layers is not sufficient to handle the complex heterogeneity of unknown watermarking algorithms in the AWPD task. Modern invisible watermarks often concentrate energy in specific narrow-band frequencies. The primary design intention of the Dynamic Multi-band Attention mechanism (DMSA) is to recalibrate channel weights in the deep feature space using multi-branch discrete cosine transforms and extremum pooling.